# Game-RL: Synthesizing Multimodal Verifiable Game Data to Boost VLMs’ General Reasoning

### v5

@@ 2.2 Game task and QA Template design
The second step is to design game task and question-answer templates (QA Templates) according to each video game. Game task designing base on the visual element and action space of video game generated in the first step. Taking Sokoban as an example, we can design a task about predicting the position of player after moves according to a game state screen, so we can propose a question and an answer. The question is "If the player moves left, up, down, down, where is the player?" and the answer is "The position of player is (2,2)." Then we can generalize these questions and answers into a QA Template, as shown in figure 1 (Step 2). Meanwhile, we can fill contents into a QA Template to get many instances. Detailed QA Templates of Sokoban is provided in Appendix 14.1.
@@ 5 Experiment
Table 4: GameQA’s generalization is competitive compared to outstanding geometry visual reasoning datasets
@@ 5.2 Main results
GameQA’s out-of-domain generalization is competitive compared to outstanding geometry visual reasoning datasets To better understand the GameQA dataset’s advantages, we compared it with the outstanding geometry visual reasoning datasets, include MAVIS [35], Multimodal-Open-R1 [16] and MultiMath [22]. MAVIS includes various geometry and function problems. Multimodal-Open-R1 is a geometry-centered dataset. MultiMath is a comprehensive and diverse multimodal mathematical dataset. Based on Qwen2.5-VL-7B, we applied the same GRPO method on 5k GameQA samples, 8k samples from MAVIS, 8k Multimodal-Open-R1 samples, 8k samples from MultiMath respectively, to conduct comparative training. The results (Table 4) show that despite having fewer training samples (5k vs. 8k) that are also out-of-domain for geometry and function tasks, the GameQA-trained model is competitive compared to its counterparts trained on geometry or function data, where general vision benchmarks would be considered in-domain. These results suggest that GameQA enables stronger out-of-domain generalization, even when using less data from a mismatched domain.
@@
Figure 3: Scaling effect of data quantity on general vision benchmarks, of Qwen2.5-VL-7B-Instruct trained on a total of 20k samples. For every 1000 samples trained, the model is tested on general vision benchmarks. We divide results into 3 stages and use bin average to show the upward trend more clearly.
Figure 4: The scaling effect of GameQA. As VLMs are trained on an increasing number of distinct games, their performance on general visual benchmarks improves.
@@ 10.3.2 Training hyperparameters
Table 8: GRPO Hyperparameters. Some hyperparameters are different when training 32B model.
@@
Learning Rate (non-32B) 2e-7
Learning Rate (32B) 5e-7
Batch Size (non-32B) 3
Batch Size (32B) 16
KL-divergence Coefficient ($\beta$) (non-32B) 0.04
KL-divergence Coefficient ($\beta$) (32B) 0.001
@@ 11.2 Training data difficulty experiment
11.3 Model scale experiment
To verify if GameQA’s benefits extend to larger models, we conducted experiments on Qwen2.5-VL-32B using LoRA. As shown in Table 13, the 32B model demonstrates greater performance gains than the 7B model under identical LoRA settings across all domains. The average improvement on general benchmarks is +0.78 for the 32B model compared to +0.68 for the 7B model. This trend suggests that GameQA is an effective resource for enhancing reasoning in larger-scale VLMs, and we hypothesize that full-parameter fine-tuning would yield even more substantial gains.
Table 13: Performance gains from GRPO (LoRA, 5k GameQA data) on 7B vs. 32B models. Gains are shown in parentheses.
Model Training General Vision In-domain Out-of-domain
Benchmarks Games Games
Qwen2.5-VL-7B Baseline 49.94 25.78 27.09
+GameQA (LoRA) 50.62 (+0.68) 30.90 (+5.12) 30.45 (+3.36)
Qwen2.5-VL-32B Baseline 60.18 35.91 35.76
+GameQA (LoRA) 60.96 (+0.78) 43.24 (+7.33) 40.77 (+5.01)

### v6

@@ 2.2 Game task and QA Template design
The second step is to design game task and question-answer templates (QA Templates) according to each video game. Game task design base on the visual element and action space of video game generated in the first step. Taking Sokoban as an example, we can design a task about predicting the position of player after moves according to a game state screen, so we can propose a question and an answer. The question is "If the player moves left, up, down, down, where is the player?" and the answer is "The position of player is (2,2)." Then we can generalize these questions and answers into a QA Template, as shown in figure 1 (Step 2). Meanwhile, we can fill contents into a QA Template to get many instances. Detailed QA Templates of Sokoban is provided in Appendix 14.1.
@@ 5 Experiment
Table 3: Evaluation results on general vision benchmarks. These results show that game-only training enhances general reasoning capability of VLMs. We fine-tuned four VLMs (InternVL2.5-8B [3], InternVL3-8B [4], Qwen2.5-VL-7B [2], and LLaVA-OneVision-7B [12]) on 5K GameQA samples using GRPO, resulting in the models Game-RL-InternVL2.5-8B, Game-RL-InternVL3-8B, Game-RL-Qwen2.5-VL-7B, and Game-RL-LLaVA-OV-7B. The percentage of performance improvements compared to the vanilla model is denoted by ($\uparrow$). Best performance per section is indicated in bold.
@@
Game-RL-LLaVA-OV-7B 42.27 (+1.04) 58.20 34.92 82.53 41.31 27.30 23.07 28.58
Table 4: GameQA’s generalization is competitive compared to outstanding geometry visual reasoning datasets. Based on Qwen2.5-VL-7B, we applied the same GRPO method on 5k GameQA samples, 8k samples from MAVIS, 8k Multimodal-Open-R1 samples, 8k samples from MultiMath respectively, a mixture of 5k GameQA and 8k MultiMath samples, to conduct comparative training.
@@ 5.2 Main results
GameQA’s out-of-domain generalization is competitive compared to outstanding geometry visual reasoning datasets To better understand the GameQA dataset’s advantages, we compared it with the outstanding geometry visual reasoning datasets, include MAVIS [35], Multimodal-Open-R1 [16] and MultiMath [22]. MAVIS includes various geometry and function problems. Multimodal-Open-R1 is a geometry-centered dataset. MultiMath is a comprehensive and diverse multimodal mathematical dataset. The results (Table 4) show that despite having fewer training samples (5k vs. 8k) that are also out-of-domain for geometry and function tasks, the GameQA-trained model is competitive compared to its counterparts trained on geometry or function data, where general vision benchmarks would be considered in-domain. These results suggest that GameQA enables stronger out-of-domain generalization, even when using less data from a mismatched domain.
@@
Figure 3: The scaling effect of training data quantity on general vision benchmarks for Qwen2.5-VL-7B-Instruct. The model was trained on a total of 20k samples (20 games) and evaluated every 1,000 samples. To clearly demonstrate the upward trend, the results are divided into three stages and presented using bin averaging (as described in Section 11.3).
Figure 4: The scaling effect of GameQA. As VLMs are trained on an increasing number of distinct games, their performance on general visual benchmarks improves. Game selection is shown in Table 7.
@@ 5.6 Advanced vision-language models perform notably worse than humans on the GameQA benchmark
Table 5: Evaluation results on GameQA benchmark. The experimental setup is consistent with Table 3: four VLMs were fine-tuned on 5K GameQA samples using GRPO. We evaluate the models on both in-domain and out-of-domain game categories. The percentage improvement over the vanilla model is marked with a green upward arrow ($\uparrow$). The best performance for each category is presented in bold.
@@ 10.3.2 Training hyperparameters
Table 8: GRPO Hyperparameters.
@@
Learning Rate 2e-7
Batch Size 3
KL-divergence Coefficient ($\beta$) 0.04